OpenStudy (anonymous):

What is a polynomial with a single root at x = -3 and a triple root at x = 4?

OpenStudy (aum):

x = -3 x + 3 = 0. Therefore, (x+3) is a factor. x = 4 x - 4 = 0. Therefore, (x-4) is a factor. And since it is a triple root, (x-4)^3 is a factor. The simplest polynomial is: (x+3)(x-4)^3. You can multiply it out if necessary.